OpenStudy (anonymous):

Find the component form of v given its magnitude and the angle it makes with the positive x-axis. ||v||=3 v in direction 3i +4j

OpenStudy (anonymous):

3i +4j has unit vector of 3/5 i + 4/5 j

OpenStudy (anonymous):

we are going 3 unit in that direction 3(3/5 i + 4/5 j)=9/5 i + 12/5 j

OpenStudy (anonymous):

finding unit vector vector/ magnitude 3 i + 4 j 3^2+ 4^2=5^2

OpenStudy (anonymous):

why is it over 5???

OpenStudy (anonymous):

we need unit vector . unit vector = vector/magnitude magnitude=5

OpenStudy (anonymous):

i don't get how you got 5

OpenStudy (anonymous):

magnitude is attained by \[a i + b j\] squaring each component , adding them together , and taking square root \[\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

ok... so i would do

OpenStudy (anonymous):